# Fully discrete error analysis of finite element discretizations of time-dependent Stokes equations in a stream-function formulation

### v1

@@ Abstract
In this paper we establish best approximation type estimates for the fully discrete Galerkin solutions of time-dependent Stokes problem using the stream-function formulation. For the time discretization we use the discontinuous Galerkin method of arbitrary degree, whereas we present the space discretization in a general framework. This makes our result applicable for a wide variety of space discretization methods, provided some Galerkin orthogonality conditions are satisfied. As an example, conformal $C^{1}$ and $C^{0}$ interior penalty methods are covered by our analysis. The results do not require any additional regularity assumptions beyond the natural regularity given by the domain and data and can be used for optimal control problems.
@@ 1 Introduction
For simply connected $\Omega$, the equations (1) and (5) are equivalent, for a proof, we refer to [17]. Although the stationary problem is well investigated, (cf. Parts 4 and 5 in [18]), it seems there is no clean theory of fully discrete finite element error estimates for Galerkin methods for the time dependent problem. For example, a relatively recent paper [1] treated backward Euler for virtual elements and derived some fully discrete error estimates in $L^{\infty}(H^{1})$ norm. However, besides treating only lowest order discretization the error analysis presented there is rather messy and involves unnecessary terms, like time derivatives for the estimation of the spacial error, as well as artificial assumptions on $\partial_{tt}\psi$ and $\mathbf{g}\in L^{\infty}(I;H^{r-1}(\Omega)^{2})$ for some $r>\frac{1}{2}$.
@@
where $\psi_{kh}$ is the fully discrete Galerkin solution for (5), $R_{h}$ is the Galerkin projection of the stationary biharmonic problem and $\pi_{k}$ is a certain time projection, and $\chi$ is an arbitrary element from the finite dimensional space. Under low regularity assumptions, i.e. $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$, using the approximation properties of the above mentioned projections we can establish the following optimal convergence result
@@ 2.1 Notation
We will use the standard notation for Lebesgue and Hilbert spaces. Most of our analysis will be $L^{2}$ based and we will denote the space and space-times inner-products by
@@
When the situation is less clear we will use the standard notation for the norms, $L^{2}(\Omega)$, $H^{2}(\Omega)$ and etc. We will denote by $\langle\cdot,\cdot\rangle$ a dual pairing between a function space and its dual.
@@ 2.2 Function spaces
$H^{1}_{0}(\Omega)$ $=\{v\in H^{1}(\Omega)\ :\ v=0\quad\text{on $\partial\Omega$}\},$
$H^{2}_{0}(\Omega)$ $=\{v\in H^{2}(\Omega)\ :\ v=\partial_{n}v=0\quad\text{on $\partial\Omega$}\},$
@@
The proof follows by the integration by parts. ∎
@@ 2.3 Results on the time-dependent Stokes problem
$\mathbf{V}:=\{\mathbf{v}\in H^{1}_{0}(\Omega)^{2}:\nabla\cdot\mathbf{v}=0\}\quad\text{ and }\quad\mathbf{H}:=\{\mathbf{v}\in L^{2}(\Omega)^{2}:\nabla\cdot\mathbf{v}=0\land\mathbf{n}\cdot\mathbf{v}=0\},$
@@
$\left\langle\partial_{t}\mathbf{u},\mathbf{v}\right\rangle_{I\times\Omega}+\left(\nabla\mathbf{u},\nabla\mathbf{v}\right)_{I\times\Omega}=\left\langle\mathbf{g},\mathbf{v}\right\rangle_{I\times\Omega}\quad\forall\mathbf{v}\in L^{2}(I;\mathbf{V}),$ (12)
where $\left\langle\cdot,\cdot\right\rangle_{I\times\Omega}$ denotes the duality pairing between $L^{2}(I;\mathbf{V})$ and $L^{2}(I;\mathbf{V}^{*})$. There holds the following result TO ADD:
@@
By [16, Lemma 3] there exists a unique $p\in H^{1}(\Omega)/\mathbb{R}$ such that $\operatorname{\mathbf{curl}}\psi_{0}-\mathbf{u}_{0}=\nabla p$. Due to the first isomorphism of (10) $\operatorname{\mathbf{curl}}\psi_{0}\in\mathbf{H}$ and by assumption $\mathbf{u}_{0}\in\mathbf{V}\subset\mathbf{H}$, it holds $\nabla p\in\mathbf{H}$. The definition of $\mathbf{H}$ then implies $\Delta p=0$ and $\partial_{\mathbf{n}}p=0$. Hence $p$ is constant and due to $p\in H^{1}(\Omega)/\mathbb{R}$ it actually vanishes. This implies $\operatorname{\mathbf{curl}}\psi_{0}=\mathbf{u}_{0}\in\mathbf{V}\subset H^{1}_{0}(\Omega)^{2}$.
By definition of $\operatorname{\mathbf{curl}}$, this implies
@@
Hence especially $\psi_{0}\in H^{2}(\Omega)$ and $\partial_{\mathbf{n}}\psi_{0}=0$ on $\partial\Omega$. This shows $\psi_{0}\in H^{2}_{0}(\Omega)$. The norm bound can then be obtained using Lemma 1 and Lemma 2, which yields
@@ 2.4 Results for stationary biharmonic equation
$w\in H^{2}_{0}(\Omega)\quad:\quad\left(\Delta w,\Delta\varphi\right)_{\Omega}=\left\langle f,\varphi\right\rangle_{\Omega}\quad\forall\varphi\in H^{2}_{0}(\Omega).$ (14)
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded Lipschitz domain and $f\in H^{-2}(\Omega)$. Then there is a unique weak solution $w\in H^{2}_{0}(\Omega)$ and there exists a constant $C$ such that
@@
$\left\langle\mathcal{A}w,\varphi\right\rangle_{\Omega}=\left(\Delta w,\Delta\varphi\right)_{\Omega}\qquad\forall\varphi\in H^{2}_{0}(\Omega)$ (17)
@@
$\left(\Delta w,\Delta\varphi\right)_{\Omega}=\left({\rm NewA}g,{\rm NewA}\varphi\right)_{\Omega}\quad\forall\varphi\in H^{2}_{0}(\Omega).$ (18)
@@
and $\left(\nabla T(g),\nabla g\right)_{\Omega}=0$ implies $T(g)=0$ and thus $g=0$. Thus, $T$ is also injective. We apply spectral theorem for compact normal operators, see, e.g., [5, Theorem 12.12], and obtain the result of the lemma. ∎
@@ 2.5 Results for time-dependent problems
$\lVert\psi\rVert_{C(\bar{I};H^{2}_{0}(\Omega))}\leq C\left(\lVert\Delta^{2}\psi\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\partial_{t}\psi\rVert_{L^{2}(I;H^{1}_{0}(\Omega))}\right).$
@@
As functions of the form of $\varphi^{N}$ are dense in $\mathbb{Y}=L^{2}(I;H^{2}_{0}(\Omega))$, we have just shown the first equation of (22). To that the initial condition is satisfied, let $\varphi\in C^{1}(\bar{I};H^{2}_{0}(\Omega))$ be arbitrary with $\varphi(T)=0$. Testing the first equation of (22) with $\varphi$, yields after integration by parts in time and space, that
@@
$=\left\langle\partial_{t}\nabla\psi_{n},\nabla P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\Delta\psi_{n},\Delta P_{n}\varphi\right)_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla P_{n}\varphi(0)\right)_{\Omega}$
$=\left\langle f,P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla P_{n}\varphi(0)\right)_{\Omega}$
@@ 2.6 Interpolated Regularity
The formulations of Theorems 11 and 13 were chosen to explicitly highlight the best regulsrity of $\psi$ that we require, but obscures the derivation of a corresponding intermediate result. Before we state the next corollary, let us note, that the stability estimate of Theorem 11 can equivalently be written as: If $f\in L^{2}(I;H^{-2}(\Omega))$ and $\psi_{0}\in H^{1}_{0}(\Omega)$, then
@@ 3.1 Spatial discretization
$\left(\nabla\Pi_{h}w,\nabla v_{h}\right)_{\Omega}=\left(\nabla w,\nabla v_{h}\right)_{\Omega}\qquad\forall v_{h}\in V_{h}.$ (40)
By definition, $\Pi_{h}$ is stable in $H^{1}_{0}(\Omega)$, i.e. there holds
@@
$a_{h}(\varphi_{h},\varphi_{h})\geq 0\quad\forall\varphi_{h}\in V_{h}\quad\text{and}\quad a_{h}(\varphi_{h},\varphi_{h})=0\quad\Rightarrow\quad\varphi_{h}=0,$
@@
$a_{h}(\phi_{h},\varphi_{h})=({\rm NewA}A_{h}\phi_{h},{\rm NewA}\varphi_{h})_{\Omega}\quad\forall\phi_{h},\varphi_{h}\in V_{h}.$ (43)
@@
$a_{h}(w_{h},\varphi_{h})=({\rm NewA}g,{\rm NewA}\varphi_{h})_{\Omega}\quad\forall\varphi_{h}\in V_{h}.$
@@
Let $\theta$ and $V_{h}$ be as in 1. For all $\psi\in X:=\mathcal{A}^{-1}(H^{-1-\theta}(\Omega))$, $\varphi_{h}\in V_{h}$, the bilinear form $a_{h}(\cdot,\cdot)$ can be extended to
$a_{h}(\cdot,\cdot):X+V_{h}\times V_{h}\to\mathbb{R},$
and there holds
@@
Under 2, we have the following form of a Galerkin orthogonality: Let the solution $\psi$ of (13) for given data $f\in H^{-1-\theta}(\Omega)$ satisfy $\psi\in X$. Then
$a_{h}(\psi-\psi_{h},\varphi_{h})=0\qquad\forall\varphi_{h}\in V_{h},$ (44)
@@
$a_{h}(\psi_{h},\varphi_{h})=\left\langle f,\varphi_{h}\right\rangle_{\Omega}\qquad\forall\varphi_{h}\in V_{h}.$
@@ 3.2 Full space-time discretization
$(\pi_{k}v-v,\phi)_{I_{m}}$ $=0\quad\forall\phi\in\mathbb{P}_{r-1}(I_{m}),\quad r>0,$ (48)
@@
where $\theta\in[0,1]$ satisfies 1. Note that due to 1, the right hand side $\left\langle f,\varphi_{kh}\right\rangle_{I\times\Omega}$ is well defined. Moreover, we have the following parabolic Galerkin orthogonality between the continuous and fully discrete solution.
@@
Due to the regularity of $f$ and $\psi_{0}$, as a consequence of the interpolated regularity result in Corollary 15, we have
@@
By 2, the regularity of $\mathcal{A}\psi$ implies $\psi\in L^{2}(I;X)$, with $X$ from Assumption 2, and thus in the definition of $B_{h}(\cdot,\cdot)$, the term $\int_{I}a_{h}(\psi,\varphi_{kh})$ is well defined. Further, as due to Theorem 11 it holds $\nabla\psi\in C(\bar{I};L^{2}(\Omega)^{2})$, all jump terms vanish and $\psi_{0}^{+}=\lim_{t\to 0+}\psi(t)=\psi(0)$. We thus obtain
@@
the regularities of $f,\partial_{t}\Delta\psi,\mathcal{A}\psi$ imply (52) and satisfy
@@
where the last step is precisely the definition of $\psi_{kh}$. This concludes the proof. ∎
@@ 4 Fully discrete error analysis
$B_{h}(\varphi_{kh},z_{kh})=({\rm NewA}(\psi_{kh}-\chi_{kh}),{\rm NewA}\varphi_{kh})_{I\times\Omega}\quad\forall\varphi_{kh}\in X_{k}^{r}(V_{h}).$ (64)
@@
where we used one dimensional inverse inequality on $I_{m}$. Combining the estimates for $J_{1}$, $J_{2}$, and $J_{3}$ and canceling $\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|_{I\times\Omega}$ on both sides, we obtain the estimate (63). From this, we obtain by inserting twice $\pm\psi$ and applying triangle inequality, that

### v2

@@ Abstract
In this paper we establish best approximation type error estimates for the fully discrete Galerkin solutions of the time-dependent Stokes problem using the stream-function formulation. For the time discretization we use the discontinuous Galerkin method of arbitrary degree, whereas we present the space discretization in a general framework. This makes our result applicable for a wide variety of space discretization methods, provided some Galerkin orthogonality conditions are satisfied. As an example, conformal $C^{1}$ and $C^{0}$ interior penalty methods are covered by our analysis. The results do not require any additional regularity assumptions beyond the natural regularity given by the domain and data and can be used for optimal control problems.
@@ 1 Introduction
For simply connected $\Omega$, the equations (1) and (5) are equivalent, for a proof, we refer to [17]. Although the stationary problem is well investigated, (cf. Parts 4 and 5 in [18]), it seems there is no clean theory of fully discrete finite element error estimates for Galerkin methods for the time dependent problem. For example, a relatively recent paper [1] treated backward Euler for virtual elements and derived some fully discrete error estimates in $L^{\infty}(H^{1})$ norm. However, in order to achieve these estimates, the authors rely on regularity in time for spacial error estimates and had to impose overly strict assumptions on $\partial_{tt}\psi$ and $\mathbf{g}\in L^{\infty}(I;H^{r-1}(\Omega)^{2})$ for some $r>\frac{1}{2}$.
@@
where $\psi_{kh}$ is the fully discrete Galerkin solution for (5), $R_{h}$ is the Galerkin projection of the stationary biharmonic problem and $\pi_{k}$ is a certain time projection, and $\chi$ is an arbitrary element from the finite dimensional space. Under low regularity assumptions, i.e. $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$, using the approximation properties of the above mentioned projections, we can establish the following optimal convergence result
@@ 2.1 Notation
We will use the standard notation for Lebesgue and Hilbert spaces. Most of our analysis will be $L^{2}$ based and we will denote the space and space-time inner-products by
@@
When the situation is less clear we will use the standard notation for the norms, $L^{2}(\Omega)$, $H^{2}(\Omega)$ etc. We will denote by $\langle\cdot,\cdot\rangle$ a dual pairing between a function space and its dual.
@@ 2.2 Function spaces
$H^{1}_{0}(\Omega)$ $=\{v\in H^{1}(\Omega)\ |\ v=0\quad\text{on $\partial\Omega$}\},$
$H^{2}_{0}(\Omega)$ $=\{v\in H^{2}(\Omega)\ |\ v=\partial_{n}v=0\quad\text{on $\partial\Omega$}\},$
@@
The proof follows by integration by parts and Poincaré inequality for the lower order terms. ∎
@@ 2.3 Results on the time-dependent Stokes problem
[…]
@@
$\left\langle\partial_{t}\mathbf{u},\mathbf{v}\right\rangle_{I\times\Omega}+\left(\nabla\mathbf{u},\nabla\mathbf{v}\right)_{I\times\Omega}=\left\langle\mathbf{g},\mathbf{v}\right\rangle_{I\times\Omega}\quad\text{for all }\mathbf{v}\in L^{2}(I;\mathbf{V}),$ (12)
where $\left\langle\cdot,\cdot\right\rangle_{I\times\Omega}$ denotes the duality pairing between $L^{2}(I;\mathbf{V})$ and $L^{2}(I;\mathbf{V}^{*})$. There holds the following result, see [30, Chapter III, Theorem 1.1 Proposition 1.2]:
@@
By [16, Lemma 3] there exists a unique $p\in H^{1}(\Omega)/\mathbb{R}$ such that $\operatorname{\mathbf{curl}}\psi_{0}-\mathbf{u}_{0}=\nabla p$. Due to the first isomorphism of (10) $\operatorname{\mathbf{curl}}\psi_{0}\in\mathbf{H}$ and by assumption $\mathbf{u}_{0}\in\mathbf{V}\subset\mathbf{H}$, it holds $\nabla p\in\mathbf{H}$. The definition of $\mathbf{H}$ then implies $\Delta p=0$ and $\partial_{\mathbf{n}}p=0$. Hence $p$ is constant and due to $p\in H^{1}(\Omega)/\mathbb{R}$ it actually vanishes. This implies $\operatorname{\mathbf{curl}}\psi_{0}=\mathbf{u}_{0}\in\mathbf{V}\subset H^{1}_{0}(\Omega)^{2}$. By definition of $\operatorname{\mathbf{curl}}$, this implies
@@
Hence especially $\psi_{0}\in H^{2}(\Omega)$ and $\partial_{\mathbf{n}}\psi_{0}=0$ on $\partial\Omega$. This shows $\psi_{0}\in H^{2}_{0}(\Omega)$. The norm bound can then be obtained using Lemmas 1 and 2, which yields
@@ 2.4 Results for stationary biharmonic equation
$w\in H^{2}_{0}(\Omega)\quad:\quad\left(\Delta w,\Delta\varphi\right)_{\Omega}=\left\langle f,\varphi\right\rangle_{\Omega}\quad\text{for all }\varphi\in H^{2}_{0}(\Omega).$ (14)
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded Lipschitz domain and $f\in H^{-2}(\Omega)$. Then (14) has a unique weak solution $w\in H^{2}_{0}(\Omega)$ and there exists a constant $C$ such that
@@
$\left\langle\mathcal{A}w,\varphi\right\rangle_{\Omega}=\left(\Delta w,\Delta\varphi\right)_{\Omega}\qquad\text{for all }\varphi\in H^{2}_{0}(\Omega)$ (17)
@@
$\left(\Delta w,\Delta\varphi\right)_{\Omega}=\left({\rm NewA}g,{\rm NewA}\varphi\right)_{\Omega}\quad\text{for all }\varphi\in H^{2}_{0}(\Omega).$ (18)
@@
and $\left(\nabla T(g),\nabla g\right)_{\Omega}=0$ implies $T(g)=0$ and thus $g=0$. Thus, $T$ is also injective. We apply the spectral theorem for compact normal operators, see, e.g., [5, Theorem 12.12], and obtain the result of the lemma. ∎
@@ 2.5 Results for time-dependent problems
Thus the space $\mathbb{X}$ defined in (22) satisfies the continuous embedding: $\mathbb{X}\hookrightarrow C(\bar{I};H^{1}_{0}(\Omega))$.
@@
$\lVert\psi\rVert_{C(\bar{I};H^{2}_{0}(\Omega))}\leq C\left(\lVert\mathcal{A}\psi\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\partial_{t}\psi\rVert_{L^{2}(I;H^{1}_{0}(\Omega))}\right).$
@@
As functions of the form of $\varphi^{N}$ are dense in $\mathbb{Y}=L^{2}(I;H^{2}_{0}(\Omega))$, we have just shown the first equation of (23). To show, that the initial condition is satisfied, let $\varphi\in C^{1}(\bar{I};H^{2}_{0}(\Omega))$ be arbitrary with $\varphi(T)=0$. Testing the first equation of (23) with $\varphi$, yields after integration by parts in time and space, that
@@
$=\left\langle\partial_{t}\nabla\psi_{n},\nabla P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\Delta\psi_{n},\Delta P_{n}\varphi\right)_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla\varphi(0)\right)_{\Omega}$
$=\left\langle f,P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla\varphi(0)\right)_{\Omega}$
@@ 2.6 Interpolated Regularity
The formulations of Theorems 11 and 13 were chosen to explicitly highlight the best regularity of $\psi$ that we require, but obscure the derivation of a corresponding intermediate result. Before we state the next corollary, let us note, that the stability estimate of Theorem 11 can equivalently be written as: If $f\in L^{2}(I;H^{-2}(\Omega))$ and $\psi_{0}\in H^{1}_{0}(\Omega)$, then
@@ 3.1 Spatial discretization
$\left(\nabla\Pi_{h}w,\nabla v_{h}\right)_{\Omega}=\left(\nabla w,\nabla v_{h}\right)_{\Omega}\qquad\text{for all }v_{h}\in V_{h}.$ (41)
By definition, $\Pi_{h}$ is stable in $H^{1}_{0}(\Omega)$, i.e., there holds
@@
$a_{h}(\varphi_{h},\varphi_{h})\geq 0\quad\text{for all }\varphi_{h}\in V_{h}\quad\text{and}\quad a_{h}(\varphi_{h},\varphi_{h})=0\quad\Rightarrow\quad\varphi_{h}=0,$
@@
$a_{h}(\phi_{h},\varphi_{h})=({\rm NewA}A_{h}\phi_{h},{\rm NewA}\varphi_{h})_{\Omega}\quad\text{for all }\phi_{h},\varphi_{h}\in V_{h}.$ (44)
@@
$a_{h}(w_{h},\varphi_{h})=({\rm NewA}g,{\rm NewA}\varphi_{h})_{\Omega}\quad\text{for all }\varphi_{h}\in V_{h}.$
@@
Let $\theta$ and $V_{h}$ be as in 1. The bilinear form $a_{h}(\cdot,\cdot)$ can be extended to
$a_{h}(\cdot,\cdot):(X+V_{h})\times(X+V_{h})\to\mathbb{R},$
where $X:=\mathcal{A}^{-1}(H^{-1-\theta}(\Omega))$. Furthermore, for all $\psi\in X,\varphi_{h}\in V_{h}$, there holds
@@
Under 2, we have the following form of a Galerkin orthogonality: Let $\psi\in H^{2}_{0}(\Omega)$ be the solution of (13) for given data $f\in H^{-1-\theta}(\Omega)$. Then $\psi\in X$, and it holds
$a_{h}(\psi-\psi_{h},\varphi_{h})=0\qquad\text{for all }\varphi_{h}\in V_{h},$ (45)
@@
$a_{h}(\psi_{h},\varphi_{h})=\left\langle f,\varphi_{h}\right\rangle_{\Omega}\qquad\text{for all }\varphi_{h}\in V_{h}.$
@@ 3.2 Full space-time discretization
$(\pi_{k}v-v,\phi)_{I_{m}}$ $=0\quad\text{for all }\phi\in\mathbb{P}_{r-1}(I_{m}),\quad r>0,$ (49)
@@
where $\theta\in(0,1]$ satisfies 1. Note that due to 1, the right hand side $\left\langle f,\varphi_{kh}\right\rangle_{I\times\Omega}$ is well defined. Moreover, we have the following parabolic Galerkin orthogonality between the continuous and fully discrete solution.
@@
Let $\varphi_{kh}\in X_{k}^{r}(V_{h})$ be arbitrary. Due to the regularity of $f$ and $\psi_{0}$, as a consequence of the interpolated regularity result in Corollary 15, we have
@@
By 2, the regularity of $\mathcal{A}\psi$ implies $\psi\in L^{2}(I;X)$, with $X$ from Assumption 2, and thus in the definition of $B_{h}(\cdot,\cdot)$, the term $\int_{I}a_{h}(\psi,\varphi_{kh})\,dt$ is well defined. Further, as due to Theorem 11 it holds $\nabla\psi\in C(\bar{I};L^{2}(\Omega)^{2})$, all jump terms vanish and $\psi_{0}^{+}=\lim_{t\to 0+}\psi(t)=\psi(0)$. We thus obtain
@@
letting $n\to\infty$, the regularities of $f$ given by assumption, as well as $\partial_{t}\Delta\psi,\mathcal{A}\psi$ from (53) yield
@@
where the last step is precisely the definition of $\psi_{kh}$. As $\varphi_{kh}$ was arbitrary, this concludes the proof. ∎
@@ 4 Fully discrete error analysis
$B_{h}(\varphi_{kh},z_{kh})=({\rm NewA}(\psi_{kh}-\chi_{kh}),{\rm NewA}\varphi_{kh})_{I\times\Omega}\quad\text{for all }\varphi_{kh}\in X_{k}^{r}(V_{h}).$ (65)
@@
where we used the one dimensional inverse inequality on $I_{m}$. Combining the estimates for $J_{1}$, $J_{2}$, and $J_{3}$ and canceling $\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|_{I\times\Omega}$ on both sides, we obtain the estimate (64). From this, we obtain by inserting twice $\pm\psi$ and applying triangle inequality, that